OpenStudy (anonymous):

3x + y = 6 x - y = 6

OpenStudy (bamboo888):

x = 3 y = -3

OpenStudy (bamboo888):

Because when you take negatives away from positives then you're adding more negatives.

OpenStudy (bamboo888):

So if x were 3 the for the first question: 3x + y = 6 3 ( x = 3) so 3x3=9 Then it would become: 9 + y = 6 So to add anything to make it a smaller number ( for positives ) then the number would have to be a negative. 9 + -? = 6. -? = -3 because 9 + -3 = 6. x - y = 6 x=3 y+-3 x(3)-y(-3)=6 ( You would know if you read my comment on negative ) "Because when you take negatives away from positives then you're adding more negatives."